Flying against the wind, a jet travels 7290 miles in 9 hours. Flying with the wind, the same jet travels 3630 miles in 3 hours. What is the rate of the jet in still air and what is the rate of the wind?
Answer by mananth(16949)

Plane speed =x mph
wind speed =y mph
against wind 9 hours
with wind 3 hours

Distance against 7290 miles distance with 3630 miles
t=d/r against wind -
7290.00 / ( x - y )= 9.00
9.00 ( x - y ) = 3.00
9.00 x - 9.00 y = 7290.00 ....................1

3630.00 / ( x + y )= 3.00
3.00 ( x + y ) = 3630.00
3.00 x + 3.00 y = 3630.00 ...............2
Multiply (1) by 1.00
Multiply (2) by 3.00
we get
9.00 x + -9.00 y = 7290.00
9.00 x + 9.00 y = 10890.00
18.00 x = 18180.00
/ 18.00
x = 1010.00 mph

plug value of x in (1) y
9.00 x -9.00 y = 7290.00
9090.00 -9.00 -9090.00 = 7290.00
-9.00 y = 7290.00
-9.00 y = -1800.00 mph
y = 200.00
Plane speed 1010 mph
wind speed 200 mph

m.ananth@hotmail.ca